# Spectrum of Schrödinger operators on subcovering graphs

### v1

@@ 2.2. Asymptotically isospectral subcoverings
ii) The triangular lattice shown in Fig. 2b is the subcovering ${\mathbb{L}}^{3}_{\mathfrak{t}_{1}}$ of the cubic lattice ${\mathbb{L}}^{3}$ with the chiral vector $\mathfrak{t}_{1}=(1,1,-1)$. It is well known that the spectrum of the Laplacian $\Delta_{\mathfrak{t}_{1}}$ on the triangular lattice has the form $\sigma(\Delta_{\mathfrak{t}_{1}})=[0,9]$. On the other hand, since $\mathfrak{t}_{1}$ is primitive and the sum of its components is odd, then, due to Example 2.10, we have $\sigma(\Delta_{\mathfrak{t}_{1}})=[0,\lambda_{1}^{+}(\mathfrak{t}_{1})]$, and the asymptotics (2.12) (which is very rude in this case) yields
@@
see Lemma 4.4. We describe all subcoverings ${\mathcal{G}}_{\mathfrak{t}}$ of the diamond lattice ${\mathcal{G}}$ isospectral to ${\mathcal{G}}$ (as well as to each other).
@@
Let ${\mathcal{G}}_{\mathfrak{t}}$ be a subcovering of the diamond lattice ${\mathcal{G}}$ with a primitive set $\mathfrak{t}\subset\Gamma$ of chiral vectors. Then ${\mathcal{G}}_{\mathfrak{t}}$ is isospectral to ${\mathcal{G}}$, i.e., $\sigma(H_{\mathfrak{t}})=\sigma(H)$ if and only if
@@
Here $H_{\mathfrak{t}}=\Delta_{\mathfrak{t}}+Q$ is the Schrödinger operator on ${\mathcal{G}}_{\mathfrak{t}}$ with the periodic potential $Q$ defined by (2.13), and $\sigma(H)$ is given by (2.14).
@@
2.5. Notions of isospectrality for periodic graphs
There are some notions of isospectrality for periodic graphs. Let $H_{1}=\Delta+Q_{1}$ and $H_{2}=\Delta+Q_{2}$ be two Schrödinger operators with periodic potentials $Q_{1}$ and $Q_{2}$ on a periodic graph ${\mathcal{G}}$.
$\bullet$ $H_{1}$ and $H_{2}$ are Floquet isospectral if
$\sigma(H_{1}(k))=\sigma(H_{2}(k))\qquad\textrm{for any}\qquad k\in{\mathcal{B}}=(-\pi,\pi]^{d},$
where $H_{s}(k)$, $s=1,2$, are the corresponding Floquet operators, see (1.4).
$\bullet$ $H_{1}$ and $H_{2}$ are Fermi isospectral if
$F_{\lambda}(H_{1})=F_{\lambda}(H_{2})\qquad\textrm{for some}\qquad\lambda\in{\mathbb{R}},$
where the Fermi surface $F_{\lambda}(H_{s})=\{k\in{\mathcal{B}}:\lambda\in\sigma(H_{s}(k))\}$, $s=1,2$.
$\bullet$ $H_{1}$ and $H_{2}$ are periodic isospectral if $\sigma(H_{1}(0))=\sigma(H_{2}(0))$, i.e., $H_{1}$ and $H_{2}$ have the same periodic spectrum.
Floquet isospectrality is the strongest one: if $H_{1}$ and $H_{2}$ are Floquet isospectral, then $H_{1}$ and $H_{2}$ are Fermi isospectral, periodic isospectral and just isospectral, i.e., $\sigma(H_{1})=\sigma(H_{2})$. The problem of Floquet isospectrality for the discrete Schrödinger operators on the lattice ${\mathbb{Z}}^{d}$ was studied in [GKT93, K89, L23]. Fermi isospectrality on ${\mathbb{Z}}^{d}$ was discussed in [L24].
Figure 4. Left: the graphs of the band functions (the dispersion relation) of the Schrödinger operator $H_{{\bf G}}$ (when $q=1$) on the hexagonal lattice ${\bf G}$. Center and right: the graphs of the band functions of the Schrödinger operators $H_{{\bf G}_{\mathfrak{t}}}$ and $H_{{\bf G}_{\mathfrak{t}_{1}}}$ on the nanotubes ${\bf G}_{\mathfrak{t}}$ and ${\bf G}_{\mathfrak{t}_{1}}$ with the chiral vectors $\mathfrak{t}=5\mathfrak{a}_{1}+2\mathfrak{a}_{2}$ (center) and $\mathfrak{t}_{1}=2\mathfrak{a}_{1}+3\mathfrak{a}_{2}$ (right). The ranges of the corresponding band functions of $H_{{\bf G}}$ and $H_{{\bf G}_{\mathfrak{t}}}$ (left and center) coincide. The ranges of the corresponding band functions of $H_{{\bf G}}$ and $H_{{\bf G}_{\mathfrak{t}_{1}}}$ (left and right) are close to each other but do not coincide.
Isospectrality of periodic graphs in the sense of coinciding their spectra as sets does not imply Floquet and even Fermi isospectrality, since it only means that the ranges of the corresponding band functions are the same, but the band functions themselves may even depend on the quasimomenta of distinct dimensions. For example, the hexagonal lattice ${\bf G}$ and the nanotube ${\bf G}_{\mathfrak{t}}$ with a primitive chiral vector $\mathfrak{t}=t_{1}\mathfrak{a}_{1}+t_{2}\mathfrak{a}_{2}$, $t_{1}-t_{2}\in 3{\mathbb{Z}}$, are isospectral ($\mathfrak{a}_{1},\mathfrak{a}_{2}$ are the periods of the hexagonal lattice), see Example 2.8. The graphs of the band functions (the dispersion relations) of the Schrödinger operators $H_{\bf G}$ on ${\bf G}$ and $H_{{\bf G}_{\mathfrak{t}}}$ on ${\bf G}_{\mathfrak{t}}$ with $\mathfrak{t}=5\mathfrak{a}_{1}+2\mathfrak{a}_{2}$ are plotted in Fig. 4 (when $q=1$). The band functions of $H_{\bf G}$ (left) depend on the two-dimensional quasimomentum $k=(k_{1},k_{2})$. The band functions of $H_{{\bf G}_{\mathfrak{t}}}$ (center) are functions of the one-dimensional quasimomentum $k$. But the ranges of these band functions are the same.
Remark 2.14.
i) For each primitive set $\mathfrak{t}\subset\Gamma$ the periodic graph ${\mathcal{G}}$ and its subcovering ${\mathcal{G}}_{\mathfrak{t}}$ have the same periodic spectrum (i.e., they are also periodic isospectral). Indeed, the Floquet operators $H(0)$ and $H_{\mathfrak{t}}(0)$ are the Schrödinger operators with the same potential $Q$ on the same fundamental graph ${\mathcal{G}}_{*}$ of ${\mathcal{G}}$ and ${\mathcal{G}}_{\mathfrak{t}}$ (see Remark 1.2 and the identity (2.1)).
ii) Notion of asymptotic isospectrality for Schrödinger operators on periodic graphs was introduced in [S24] when studying the behavior of the spectrum of the Schrödinger operators on periodic graphs under perturbations of graphs by adding edges (between existing vertices) in a periodic way.
@@ 3. Proof of the main results
The dispersion relation for the Schrödinger operator $H_{\bf G}=\Delta_{\bf G}+Q$ on ${\bf G}$ with a $\Gamma$-periodic potential $Q$ defined by (2.8) is plotted in Fig. 4 (left), when $q=1$. The Brillouin zone is parameterized by the quasimomentum $k$. The band functions of the Schrödinger operator $H_{{\bf G}_{\mathfrak{t}_{1}}}$ on the nanotube ${\bf G}_{\mathfrak{t}_{1}}$ are the restrictions of the dispersion relation for $H_{\bf G}$ to the segment of the line $L_{0}$ given by the equation $02k_{1}+3k_{2}=$ 0 (or, $\kappa_{1}=0$, in terms of $\kappa$) inside the Brillouin zone $\widetilde{\mathcal{B}}$, see Fig. 4 (right).
@@
3.2. Asymptotically isospectral graphs.
@@
We prove Propositions 2.2 and 2.4 about asymptotic isospectrality and just isospectrality of periodic graphs and their subcoverings.
@@
where $\lambda_{j}(k)$ are the band functions of the Schrödinger operator $H$ on the periodic graph ${\mathcal{G}}$, and the set ${\mathcal{B}}_{\mathfrak{t}}\subset{\mathcal{B}}:=(-\pi,\pi]^{d}$ is defined by (1.10). Combining (3.7) and (3.5), we obtain that the spectrum of $H_{\mathfrak{t}}$ has the form (2.2), where the last two inequalities follow from the inclusion ${\mathcal{B}}_{\mathfrak{t}}\subset{\mathcal{B}}$.
Now we prove the identities (2.3). Let $j\in{\mathbb{N}}_{\nu}$. We prove (2.3) for $\lambda_{j}^{-}(\mathfrak{t})$. The proof for $\lambda_{j}^{+}(\mathfrak{t})$ is similar. Due to $2\pi{\mathbb{Z}}^{d}$-periodicity of the band function $\lambda_{j}(k)$, we have
@@
where $\tau$ is the smallest eigenvalue of $({\mathcal{T}}{\mathcal{T}}^{\top})^{1/2}$. Using the first inequality in (2.2) and $2\pi{\mathbb{Z}}^{d}$-periodicity of the band function $\lambda_{j}(k)$, we have
@@
Proof of Proposition 2.4. We prove this proposition for the band edge $\lambda_{j}^{-}(\mathfrak{t})$. The proof for $\lambda_{j}^{+}(\mathfrak{t})$ is similar. Let the condition (2.5) be fulfilled for some $k_{o}\in K_{j}^{-}$. Then $k_{o}\in{\mathcal{B}}_{\mathfrak{t}}$ (see (1.10)) and,
This and the first inequality in (2.2) yield $\lambda_{j}^{-}(\mathfrak{t})=\lambda_{j}^{-}$.
i) Let $0\in K_{j}^{-}$. For $k_{o}=0$ the condition (2.5) holds. Then $\lambda_{j}^{-}(\mathfrak{t})=\lambda_{j}^{-}$.
ii) Let $k_{o}:=(\pi,\ldots,\pi)\in K_{j}^{-}$ and the sum of the entries in each row of ${\mathcal{T}}$ be even. Then the condition (2.5) holds, and $\lambda_{j}^{-}(\mathfrak{t})=\lambda_{j}^{-}$.  $\Box$
@@
Proof of Theorem 2.6. We prove (2.6) for $\lambda_{j}^{-}(\mathfrak{t})$. The proof for $\lambda_{j}^{+}(\mathfrak{t})$ is similar. Due to (3.5) and $2\pi{\mathbb{Z}}^{d}$-periodicity of the band function $\lambda_{j}(k)$, we have
@@ 4. Examples
In this section we consider simple examples of periodic graphs and describe the spectrum of the Schrödinger operators on their subcoverings.
@@ 4.1. Nanotubes
The hexagonal lattice ${\bf G}$ is a $\Gamma$-periodic graph, where $\Gamma\subset{\mathbb{R}}^{2}$ is the lattice with the basis $\mathfrak{a}_{1},\mathfrak{a}_{2}$, see Fig. 1a. The fundamental graph ${\bf G}_{*}$ of ${\bf G}$ consists of two vertices $v_{1},v_{2}$ and three multiple edges connecting these vertices, see Fig. 1d. First we describe the spectrum of the Schrödinger operator $H=\Delta+Q$ on ${\bf G}$ with a $\Gamma$-periodic potential $Q$.
Lemma 4.1.
The spectrum of the Schrödinger operator $H=\Delta+Q$ on the hexagonal lattice ${\bf G}$ with the potential $Q$ defined by (2.8) is given by
@@
Proof. For the hexagonal lattice ${\bf G}$, the Floquet operator $H(k)$ defined by (1.4) in the standard basis of $\ell^{2}(\{v_{1},v_{2}\})\cong{\mathbb{C}}^{2}$ has the following matrix:
The eigenvalues of $H(k)$ are given by
We have
$\min\limits_{k\in{\mathcal{B}}}f(k)=f\big{(}\pm\frac{2\pi}{3}\,,\mp\frac{2\pi}{3}\big{)}=0,\qquad\max\limits_{k\in{\mathcal{B}}}f(k)=f(0,0)=9.$
Then, using the definition of the band edges $\lambda_{j}^{\pm}$ in (1.6), we obtain that the spectrum of $H$ has the form (4.1), (4.2).  $\Box$
@@
i) Let $t_{1}-t_{2}\in 3{\mathbb{Z}}$. Then for the points $k_{o}=\big{(}\pm\frac{2\pi}{3}\,,\mp\frac{2\pi}{3}\big{)}\in K_{1}^{+}=K_{2}^{-}$ the condition (2.5) is fulfilled. Thus, due to Proposition 2.4 and the identities (4.1),
@@
In the following example we consider a nanotube ${\bf G}_{\mathfrak{t}}$ with a non-primitive vector $\mathfrak{t}$.
Example 4.2.
Let ${\bf G}_{\mathfrak{t}}$ be the zig-zag nanotube with $\mathfrak{t}=2\mathfrak{a}_{1}$, see Fig. 1c. Then the spectrum $\sigma(H_{\mathfrak{t}})$ of the Schrödinger operator $H_{\mathfrak{t}}=\Delta_{\mathfrak{t}}+Q$ on ${\bf G}_{\mathfrak{t}}$ with the periodic potential $Q$ defined by (2.8) has the form
$\begin{array}[]{l}\sigma(H_{\mathfrak{t}})=\sigma_{ac}(H_{\mathfrak{t}})\cup\sigma_{fb}(H_{\mathfrak{t}}),\qquad\sigma_{fb}(H_{\mathfrak{t}})=\{3\pm\sqrt{1+q^{2}}\,\},\\[8.0pt] \sigma_{ac}(H_{\mathfrak{t}})=\big{[}3-\sqrt{9+q^{2}}\,,3-\sqrt{1+q^{2}}\,\big{]}\cup\big{[}3+\sqrt{1+q^{2}}\,,3+\sqrt{9+q^{2}}\,\big{]}.\end{array}$ (4.6)
Proof. The zigzag nanotube ${\bf G}_{\mathfrak{t}}$, $\mathfrak{t}=2\mathfrak{a}_{1}$, is a periodic graph with period $\mathfrak{a}_{2}$, see Fig. 1c. The fundamental graph $\widetilde{\bf G}_{*}=({\mathcal{V}}_{*},{\mathcal{E}}_{*})$ of ${\bf G}_{\mathfrak{t}}$ is shown in Fig. 1e, where ${\mathcal{V}}_{*}=\{v_{1},v_{2},v_{3},v_{4}\}$. For the nanotube ${\bf G}_{\mathfrak{t}}$, the matrix of the Floquet operator $H_{\mathfrak{t}}(k)$, $k\in(-\pi,\pi]$, defined by (1.4) in the standard basis of $\ell^{2}({\mathcal{V}}_{*})\cong{\mathbb{C}}^{4}$ has the form
$H_{\mathfrak{t}}(k)=\left(\begin{array}[]{cc}A(k)&B\\ B&A(k)\end{array}\right),\quad A(k)=\left(\begin{array}[]{cc}3+q&-1-e^{ik}\\ -1-e^{-ik}&3-q\end{array}\right),\quad B=\left(\begin{array}[]{cc}0&-1\\ -1&0\end{array}\right).$
The eigenvalues of $H_{\mathfrak{t}}(k)$ are given by
$\lambda_{1,4}(k)=3\pm\sqrt{5+4\cos k+q^{2}}\,,\qquad\lambda_{2,3}=3\pm\sqrt{1+q^{2}}\,.$
Then, using the definition of the band edges $\lambda_{j}^{\pm}$ in (1.6), we obtain that the spectrum of $H_{\mathfrak{t}}$ has the form (4.6).  $\Box$
Remark 4.3.
The spectra of the Schrödinger operators $H=\Delta+Q$ on the hexagonal lattice ${\bf G}$ and $H_{\mathfrak{t}}=\Delta_{\mathfrak{t}}+Q$ on any nanotube ${\bf G}_{\mathfrak{t}}$ with a primitive chiral vector $\mathfrak{t}$ (the periodic potential $Q$ is defined by (2.8)) consist of two spectral bands (see Example 2.8). The spectrum of the Schrödinger operator $H_{p\mathfrak{t}}=\Delta_{p\mathfrak{t}}+Q$ on a nanotube ${\bf G}_{p\mathfrak{t}}$ with the non-primitive chiral vector $p\mathfrak{t}$, where integer $p\geqslant 2$, has $2p$ bands. For example, the spectrum of $H_{2\mathfrak{a}_{1}}$ on the zigzag nanotube ${\bf G}_{2\mathfrak{a}_{1}}$ consists of four spectral bands, two of which are flat bands, see (4.6). But, due to (1.12), the inclusion $\sigma(H_{p\mathfrak{t}})\subseteq\sigma(H)$ still holds.

### v2

@@ 2.2. Spectra of Schrödinger operators on subcoverings with primitive sets of chiral vectors
ii) The triangular lattice shown in Fig. 2b is the subcovering ${\mathbb{L}}^{3}_{\mathfrak{t}_{1}}$ of the cubic lattice ${\mathbb{L}}^{3}$ with the chiral vector $\mathfrak{t}_{1}=(1,1,-1)$. It is known that the spectrum of the Laplacian $\Delta_{\mathfrak{t}_{1}}$ on the triangular lattice has the form $\sigma(\Delta_{\mathfrak{t}_{1}})=[0,9]$. On the other hand, since $\mathfrak{t}_{1}$ is primitive and the sum of its components is odd, then, due to Example 2.8, we have $\sigma(\Delta_{\mathfrak{t}_{1}})=[0,\lambda_{1}^{+}(\mathfrak{t}_{1})]$, and the asymptotics (2.11) (which is very rude in this case) yields
@@
see, e.g., Lemma 4.2 in [S24]. We describe all subcoverings ${\mathcal{G}}_{\mathfrak{t}}$ of the diamond lattice ${\mathcal{G}}$ isospectral to ${\mathcal{G}}$ (as well as to each other).
@@
Let ${\mathcal{G}}_{\mathfrak{t}}$ be a subcovering of the diamond lattice ${\mathcal{G}}$ with a primitive set $\mathfrak{t}\subset\Gamma$ of chiral vectors, and let $H_{\mathfrak{t}}=\Delta_{\mathfrak{t}}+Q$ be the Schrödinger operator on ${\mathcal{G}}_{\mathfrak{t}}$ with the periodic potential $Q$ defined by (2.12). Then $\sigma(H_{\mathfrak{t}})=\sigma(H)$ if and only if
@@
Here $\sigma(H)$ is given by (2.13).
@@ 3. Proof of the main results
The dispersion relation for the Schrödinger operator $H_{\bf G}=\Delta_{\bf G}+Q$ on ${\bf G}$ with a $\Gamma$-periodic potential $Q$ defined by (2.7) is plotted in Fig. 5 (left), when $q=1$. The Brillouin zone is parameterized by the quasimomentum $k$. The band functions of the Schrödinger operator $H_{{\bf G}_{\mathfrak{t}_{1}}}$ on the nanotube ${\bf G}_{\mathfrak{t}_{1}}$ are the restrictions of the dispersion relation for $H_{\bf G}$ to the segment of the line $L_{0}$ given by the equation $02k_{1}+3k_{2}=$ 0 (or, $\kappa_{1}=0$, in terms of $\kappa$) inside the Brillouin zone $\widetilde{\mathcal{B}}$, see Fig. 5 (right). The ranges of the band functions of $H_{{\bf G}}$ and $H_{{\bf G}_{\mathfrak{t}_{1}}}$ (i.e., the spectra of $H_{{\bf G}}$ and $H_{{\bf G}_{\mathfrak{t}_{1}}}$) have the form
$\sigma(H_{{\bf G}})=[3-\sqrt{10}\,,2]\cup[4,3+\sqrt{10}\,],\qquad\sigma(H_{{\bf G}_{\mathfrak{t}_{1}}})\approx[3-\sqrt{10}\,,1{.}92]\cup[4{.}08,3+\sqrt{10}\,].$
They are close to each other but do not coincide.
Figure 5. Left: the graphs of the band functions (the dispersion relation) of the Schrödinger operator $H_{{\bf G}}$ (when $q=1$) on the hexagonal lattice ${\bf G}$. Right: the graphs of the band functions of the Schrödinger operator $H_{{\bf G}_{\mathfrak{t}_{1}}}$ on the nanotube ${\bf G}_{\mathfrak{t}_{1}}$ with the chiral vector $\mathfrak{t}_{1}=2\mathfrak{a}_{1}+3\mathfrak{a}_{2}$. The ranges of the corresponding band functions of $H_{{\bf G}}$ and $H_{{\bf G}_{\mathfrak{t}_{1}}}$ are close to each other but do not coincide.
@@
3.2. Spectra of Schrödinger operators on subcoverings with primitive sets of chiral vectors.
@@
We prove Proposition 2.2 about spectra of Schrödinger operators on subcovering graphs ${\mathcal{G}}_{\mathfrak{t}}$ with primitive sets $\mathfrak{t}$ of chiral vectors.
@@
where $\lambda_{j}(k)$ are the band functions of the Schrödinger operator $H$ on the periodic graph ${\mathcal{G}}$, and the set ${\mathcal{B}}_{\mathfrak{t}}\subset{\mathcal{B}}:=(-\pi,\pi]^{d}$ is defined by (1.10). Combining (3.7) and (3.5), we obtain that the spectrum of $H_{\mathfrak{t}}$ has the form (2.2), where the last inclusion follows from the definition (1.6) of the band edges $\lambda_{j}^{\pm}$ and the inclusion ${\mathcal{B}}_{\mathfrak{t}}\subset{\mathcal{B}}$.
i) We prove this item for the band edge $\lambda_{j}^{-}(\mathfrak{t})$. The proof for $\lambda_{j}^{+}(\mathfrak{t})$ is similar. Let the condition (2.3) be fulfilled for some $k_{o}\in K_{j}^{-}$. Then $k_{o}\in{\mathcal{B}}_{\mathfrak{t}}$ (see (1.10)) and,
This and the last inclusion in (2.2) yield $\lambda_{j}^{-}(\mathfrak{t})=\lambda_{j}^{-}$.
Let $0\in K_{j}^{-}$. For $k_{o}=0$ the condition (2.3) holds. Then $\lambda_{j}^{-}(\mathfrak{t})=\lambda_{j}^{-}$.
Let $k_{o}:=(\pi,\ldots,\pi)\in K_{j}^{-}$ and the sum of the entries in each row of ${\mathcal{T}}$ be even. Then the condition (2.3) holds, and $\lambda_{j}^{-}(\mathfrak{t})=\lambda_{j}^{-}$.
ii) Let $j\in{\mathbb{N}}_{\nu}$. We prove the identity (2.4) for $\lambda_{j}^{-}(\mathfrak{t})$. The proof for $\lambda_{j}^{+}(\mathfrak{t})$ is similar. Due to $2\pi{\mathbb{Z}}^{d}$-periodicity of the band function $\lambda_{j}(k)$, we have
@@
where $\tau$ is the smallest eigenvalue of $({\mathcal{T}}{\mathcal{T}}^{\top})^{1/2}$. Using the last inclusion in (2.2) and $2\pi{\mathbb{Z}}^{d}$-periodicity of the band function $\lambda_{j}(k)$, we have
@@
Proof of Theorem 2.4. We prove the asymptotics (2.5) for $\lambda_{j}^{-}(\mathfrak{t})$. The proof for $\lambda_{j}^{+}(\mathfrak{t})$ is similar. Due to (3.5) and $2\pi{\mathbb{Z}}^{d}$-periodicity of the band function $\lambda_{j}(k)$, we have
@@ 4. Examples
In this section we consider simple examples of periodic graphs and describe the spectra of the Schrödinger operators on their subcoverings.
@@ 4.1. Nanotubes
The hexagonal lattice ${\bf G}$ is a $\Gamma$-periodic graph, where $\Gamma\subset{\mathbb{R}}^{2}$ is the lattice with the basis $\mathfrak{a}_{1},\mathfrak{a}_{2}$, see Fig. 1a. The fundamental graph ${\bf G}_{*}$ of ${\bf G}$ consists of two vertices $v_{1},v_{2}$ and three multiple edges connecting these vertices, see Fig. 1d.
First we briefly describe the spectrum of the Schrödinger operator $H=\Delta+Q$ on ${\bf G}$ with a $\Gamma$-periodic potential $Q$ defined by (2.7), for more details see, e.g., [KP07]. The Floquet operator $H(k)$ given by (1.4) in the standard basis of $\ell^{2}(\{v_{1},v_{2}\})\cong{\mathbb{C}}^{2}$ has the following matrix:
The eigenvalues of $H(k)$ are given by
Then, using the definition of the band edges $\lambda_{j}^{\pm}$ in (1.6), we obtain that the spectrum of the Schrödinger operator $H=\Delta+Q$ on the hexagonal lattice ${\bf G}$ is given by
@@
i) Let $t_{1}-t_{2}\in 3{\mathbb{Z}}$. Then for the points $k_{o}=\big{(}\pm\frac{2\pi}{3}\,,\mp\frac{2\pi}{3}\big{)}\in K_{1}^{+}=K_{2}^{-}$ the condition (2.3) is fulfilled. Thus, due to Proposition 2.2.i and the identities (4.2),